OpenStudy (anonymous):

find inverse f(x) = 3x/(x-6)

OpenStudy (p0sitr0n):

take f(x)=y y=3x/(x-6) y(3x-6)=3x 3xy-6y=3x 3xy-3x=6y x=2y/(y-1)

OpenStudy (p0sitr0n):

also note that f(6) is undefined

OpenStudy (anonymous):

it is saying the answer is incorrect

OpenStudy (campbell_st):

write it as \[y = \frac{3x}{x - 6}\] and the domain restriction is \[x \neq 6\] swap x and y \[x = \frac{3y}{y - 6}\] now make y the subject of the equation.

OpenStudy (campbell_st):

multiply both sides by y - 6 \[x(y - 6) = 3y\] distribute \[xy - 6x = 3y\] subtract xy from both sides of the equation \[-6x = 3y - xy\] factor the left hand side \[-6x = y(3 - x)\] I'll leave you to finish it off..